The area of circle B is 121 times greater than the area of circle A. The radius of circle B is 44. What is the radius of circle A?

Answer by stanbon(75887)

The area of circle B is 121 times greater than the area of circle A. The radius of circle B is 44. What is the radius of circle A?
-------
Area of A = pi*r^2
Area of B = pi*44^2
---------
Equation:
pi*44^2 = 121*pi*r^2
------
121r^2 = 44^2
---
r^2 = 44^2/11^2
----
r^2 = 4^2
r = 4
----------------